# Adapting Multilingual Embedding Models to Historical Luxembourgish

### v1

@@ 5 Conclusions
In this work, we present a case study on embedding imperfectly digitized historical Luxembourgish texts. As off-the-shelf models prove suboptimal for this task, we propose a method to synthesize reliable data for adapting embedding models to this domain. We generate parallel sentence-segmented documents by translating historical Luxembourgish newspaper articles into two related higher-resource languages (German and French) using GPT-4o as a machine translation system.
To establish a hold-out test set, we reserve 233 articles and frame a bitext mining task for retrieving historical Luxembourgish sentences. Evaluating various multilingual embedding models on this task, we demonstrate that our adaptation approach improves performance by up to 13 percentage points. While our adapted models perform effectively across different historical periods, this improvement comes at the cost of reduced performance in natural language understanding (NLU) for clean contemporary texts. This trade-off highlights the need for future research into embedding historical digitized texts affected by significant OCR errors.
@@
Our findings present both positive and negative aspects. On the one hand, our method demonstrates clear benefits for the targeted use case, effectively embedding heterogeneous digitized historical texts and revealing shortcomings in off-the-shelf models. On the other hand, performance drops in control tasks on contemporary texts raise concerns. Does improved robustness to digitization errors come at the cost of weaker language understanding and reduced sensitivity to semantic nuances? Can future training strategies mitigate this trade-off? These important questions remain open for future research.
Acknowledgements
This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
@@ Appendix A Appendix
| Newspaper | Year | Sentence | LB Comprehensibility Rating | DE Translation Rating | FR Translation Rating |
@@
DE: Sie besteht aus 18.000 Röhren, wiegt 30 Tonnen und kostet 400.000 Dollar. FR: Elle est composée de 18 000 tubes, pèse 30 tonnes et coûte 400 000 dollars. |
Comprehensible | Adequate | Adequate |
@@
DE: „Das nicht, aber ich habe keinen sauberen Teller mehr. FR: "Certainement pas, mais je n’ai plus d’assiette propre," |
Comprehensible | Adequate | Adequate |
@@
DE: Der Batti: Jetzt hätten wir die Stimmung gut geschafft, hätten wir jetzt auch noch Regen. FR: Le Batti : Maintenant, nous aurions bien passé l’ambiance, si seulement nous avions aussi encore Rén. |
Confidently Guessable | Adequate | Adequate |
@@
DE: Zur Melodie: Das ist im Leben hässlich eingerichtet. FR: À la manière de : Cela est arrangé vilaine dans la vie. |
Confidently Guessable | Adequate | Adequate |
@@
DE: Wenn auch die Liss’ché weiß zu feilschen. FR: Même si Liss’ché sait lutter. |
Incomprehensible | / | / |
@@
DE: In vielen Dörfern sind alle krank, so dass die Felder nicht bestellt werden können. FR: Dans de nombreux villages, tous sont malades, si bien que les champs ne peuvent pas être cultivés. |
Incomprehensible | / | / |

### v3

@@ 5 Conclusions
In this work, we explore the adaptation of multilingual embedding models to digitized historical LB texts, a task where off-the-shelf models struggle due to limited exposure and their reliance on clean modern data. To address this issue, we generate parallel sentence-segmented documents by translating historical Luxembourgish newspaper articles into French, English, and German using GPT-4o.
To evaluate the effectiveness of adaptation, we design a historical bitext mining task with a held-out test set of 233 articles. Our results show that adaptation to parallel historical data improves retrieval accuracy by up to 13pp. However, this adaptation introduces trade-offs, particularly reducing performance on modern LB tasks that require high semantic precision, such as adversarial paraphrase detection. We mitigate this problem through a balanced data mixing strategy that helps preserve modern LB performance while improving historical text semantic search capabilities.
These results demonstrate the effectiveness of domain adaptation for historical text processing and suggest that such approaches could benefit low-resource languages facing digitization challenges. Such improvements are particularly relevant for libraries and archives, where effective cross-lingual semantic search can improve the discoverability of historical documents and support digital exploration.
@@
Our findings paves the way for better semantic search systems within Luxembourgish archives. On the one hand, our method demonstrates clear benefits for the targeted use case, effectively embedding heterogeneous digitized historical texts and revealing shortcomings in off-the-shelf models. Through our exploration of adaptation methodologies, we have produced practical embeddings for semantic search while mostly preserving modern LB performance. On the other hand, we have not strictly reached a single best model across all evaluation sets. For example, in all of our adapted models, performance on ParaLux declines, possibly indicating interference with modern LB understanding and reduced sensitivity to semantic nuances.
Overall, we have applied a single adaptation method for each model type across all available data mixes, ensuring alignment with the models’ initial training methods. Exploring alternative adaptation approaches may reveal additional patterns.
One problem with our evaluation is that they are all at the sentence level, whereas applications of such models would often be at the paragraph, article, or document level. The hypothesis that our improved performance would be reflected when embedding longer segments of text is possible, but not guaranteed. Lastly, while our research focuses on historical Luxembourgish, our methodology may also be useful for developing semantic search models in other underrepresented languages, which we do not examine in this study.
Acknowledgments
We would like to thank Fred Phillipy for helping with the human annotation. This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
@@ Appendix A Appendix
| Newspaper | Year | Sentence | LB Compr. | FR Compr. | EN Compr. | DE Compr. |
@@
FR: Elle est composée de 18 000 tubes, pèse 30 tonnes et coûte 400 000 dollars. EN: It consists of 18,000 tubes, weighs 30 tons, and costs 400,000 dollars. DE: Sie besteht aus 18.000 Röhren, wiegt 30 Tonnen und kostet 400.000 Dollar. |
Comprehensible | Adequate | Adequate | Adequate |
@@
FR: "Certainement pas, mais je n’ai plus d’assiette propre," EN: "Not really, but I don’t have a single clean plate anymore." DE: „Das nicht, aber ich habe keinen sauberen Teller mehr. |
Comprehensible | Adequate | Adequate | Adequate |
@@
FR: Le Batti : Maintenant, nous aurions bien passé l’ambiance, si seulement nous avions aussi encore Rén. EN: Batti: Now we would have a good atmosphere if we also had some rain. DE: Der Batti: Jetzt hätten wir die Stimmung gut geschafft, hätten wir jetzt auch noch Regen. |
Confidently Guessable | Adequate | Adequate | Adequate |
@@
FR: À la manière de : Cela est arrangé vilaine dans la vie. EN: To the tune: It is poorly arranged in life. DE: Zur Melodie: Das ist im Leben hässlich eingerichtet. |
Confidently Guessable | Adequate | Adequate | Adequate |
@@
FR: Même si Liss’ché sait lutter. EN: Even if Lisette knows how to flirt. DE: Wenn auch die Liss’ché weiß zu feilschen. |
Incomprehensible | / | / | / |
@@
FR: Dans de nombreux villages, tous sont malades, si bien que les champs ne peuvent pas être cultivés. EN: In many places, all are sick, so that the fields cannot be tended. DE: In vielen Dörfern sind alle krank, so dass die Felder nicht bestellt werden können. |
Incomprehensible | / | / | / |